OpenStudy (anonymous):

find the equation of the line useing the following info show all your work (-3,2),m=2

OpenStudy (anonymous):

this is so easy. just use the formula.

OpenStudy (anonymous):

(y-ysub1)=m(x-xsub1) (y-2)=2(x-(-3)) y-2=2(x+3) y-2=2x+6 2x-y=-2-6 2x-y=-8 This is the equation.

OpenStudy (anonymous):

you can check it.